OpenStudy (anonymous):

[Calculus 2] Solve the integral (question in comments).

OpenStudy (anonymous):

If someone could explain this to me, it would be great :)

OpenStudy (anonymous):

I believe there's a general reduction formula for this?

OpenStudy (anonymous):

@Shadowys What is it?

hartnn (hartnn):

hint : write sec^4 x = (sec^2 x )^2 = (1+tan^2 x)^2 then u have powers of tan x.

OpenStudy (anonymous):

"hint : write sec^4 x = (sec^2 x )^2 = (1+tan^2 x)^2 then u have powers of tan x." Wow...I feel stupid... xD

OpenStudy (anonymous):

Yeah, that works. Thanks! :)

hartnn (hartnn):

welcome ^_^